OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

x = # of human; y = # of horses x + y = 74 2x + 4y = 196 x = 98 - 2y y = 98-74 = 28 x = 46 46 humans and 28 horses

OpenStudy (anonymous):

but the correct ans is: 24 horses and 50 humans.

OpenStudy (anonymous):

yeah my bad i put 98-74=28 and it should be equal to 24. that would give your answer

OpenStudy (anonymous):

yeah thanks... got it